# New Methods for Network Count Time Series

### v1

@@ 2.2.1 NGNAR model definition
  • The exponential response: $g(x)=\exp(x)$: in which case the model is similar to the log-linear Poisson autoregression model in Fokianos and Tjøstheim (2011). The exponential response clearly assumes a linear relationship under log-transformation of the target, which is not always the case and needs to be checked.
@@
Our exposition below mainly uses the $\operatorname{softplus}$ response function with $c=1$, i.e., $s_{c}(x)=s_{1}(x)$. Estimation can be performed either by conditional least squares or conditional maximum likelihood and both are discussed below.
@@ 2.2.3 Remarks on conditional least squares estimation
Again, for notational simplicity, we consider the global $\alpha$ case. Let $\mathbf{\beta}=(\alpha_{1},\beta_{1,1},\dots,\beta_{1,s_{1}},\dots,\alpha_{p},\beta_{p,1},\dots,\beta_{p,s_{p}},\alpha_{0})^{T}$ be the parameter of interest, $\mathbf{y}$ be the target vector as defined in (26), and $X$ be the design matrix as defined in (27). Then, the conditional least squares estimator, $\hat{\mathbf{\beta}}$, is the one that minimizes $||Y-\mathbf{g}(X\mathbf{\beta})||^{2}$, where $\mathbf{g}(\mathbf{x})=\{g(x_{1}),g(x_{2}),\ldots\}^{T}$.
As the process is no longer linear by design, we can not use the usual linear least squares estimation method. Instead, we can use numerical methods such as gradient descent or the ADAM optimization in Kingma and Ba (2015). Choosing the solution to $X^{T}X\beta=X^{T}Y$ as the initial value can speed up the optimization. The asymptotic properties discovered for the conditional least squares estimator above for the GNARI model hold here as well.
2.2.4 Conditional maximum likelihood estimation
Unlike the GNARI model, the NGNAR model explicitly defines the conditional distribution of $X_{i,t}|\mathcal{F}_{t-1}$. Thus, it is feasible to implement the conditional maximum likelihood estimator (MLE). The conditional MLE estimator $\hat{\mathbf{\beta}}_{M}$ maximizes the conditional likelihood
$L(\mathbf{\beta})=\prod_{t=p}^{n-1}\prod_{i=1}^{N}f(\mathbf{X_{i,t+1}}|\mathcal{F}_{t},\mathbf{\beta}),$ (35)
where $f(\mathbf{X_{i,t+1}}|\mathcal{F}_{t},\mathbf{\beta})$ is the density function for $X_{i,t+1}|\mathcal{F}_{t}$. We also have that $\mathbb{E}[\mathbf{X_{i,t+1}}|\mathcal{F}_{t},\mathbf{\beta}]=g([X]_{t-p+i}\cdot\mathbf{\beta})$, where $[X]_{r}$ is the $r^{th}$ row of matrix $X$ defined in (27). We can then recognize that the objective function under the NGNAR model assumption is of the same form as that of a generalized linear model (GLM). Thus, the solution for $\hat{\mathbf{\beta}}_{M}$ can be computed using the same techniques used for GLM such as iterative weighted least squares.
If we assumed that the conditional distribution was Poisson, then the parameters estimated by conditional least squares would differ from those estimated by conditional MLE unless $X_{i,t}$ is large enough so that the Poisson conditional distribution can be approximated with a Gaussian conditional distribution. Then the two estimation methods should give similar results.
@@ 3.1 GNARI parameter estimation simulation study
This section investigates estimation performance using conditional least squares for a Poisson-GNARI(1,[1]1delimited-[]11,[1]1 , [ 1 ]) process with $\alpha_{1}=0.5$, $\beta_{1,1}=0.4$, and $\lambda=10$ on a static five vertex network shown with its adjacency matrix in Figure 1.
$M=\begin{bmatrix}0&1&0&1&1\\ 1&0&0&0&0\\ 0&0&0&1&1\\ 0&0&1&0&1\\ 1&0&1&0&0\end{bmatrix}$ (37)
Figure 1: Five-node network for simulation experiments. Left: picture of the network. Right: adjacency matrix.
We will simulate realizations from the Poisson-GNARI(1,[1]1delimited-[]11,[1]1 , [ 1 ]) process with lengths $T=50,200$ and $500$ observations and repeat this $1000$ times for each choice of $T$ and estimate parameters for each realization.
@@
50 0.482 (0.053) 0.370 (0.061) 14.8 (6.57)
200 0.496 (0.025) 0.392 (0.031) 11.3 (2.96)
500 0.497 (0.015) 0.397 (0.019) 10.6 (1.73)
[…]
Table 1: The mean (and standard deviation) over $1000$ conditional least squares estimates of each parameter in the GNARI(1,[1]1delimited-[]11,[1]1 , [ 1 ]) model for each length $T$
Table 1 shows the results: the mean of the estimates clearly approaches the truth as $T$ gets larger. The standard deviation is approximately inversely proportional to $\sqrt{T}$, i,e., $\frac{\text{std}(\hat{\theta}_{T})}{\text{std}(\hat{\theta}_{m})}\approx\frac{m}{T}$, which is consistent with the asymptotic properties. It is also worth noting that there is a tendency for underestimation of $\alpha_{1}$ and $\beta_{1,1}$, but overestimation of $\lambda$.
@@ 3.2 NGNAR parameter estimation simulation study
Table 2 shows the results of a similar simulation study to the previous one, but for a NGNAR process. We simulate one hundred realizations of a Poisson-NGNAR(1,[1]1delimited-[]11,[1]1 , [ 1 ]) with $g(\cdot)=\operatorname{softplus}(\cdot)$, $\alpha_{1}=0.5$, $\beta_{1,1}=-0.4$, and $\alpha_{0}=10$ for lengths $T=50,200,500$ on the same network as in the previous section. We only use $100$ realizations as NGNAR estimation is computationally intensive due to its nonlinear nature.
For each realization, we fit the NGNAR model using both conditional least squares and conditional maximum likelihood with optimisation by the ADAM optimiser from Kingma and Ba (2015) and both under TensorFlow as described in Abadi et al. (2015).
@@
[…]
Table 2: The mean (and standard deviation) across the 100 conditional least squares and conditional MLE estimates for each parameter in the Poisson-NGNAR(1,[1]1delimited-[]11,[1]1 , [ 1 ]) model for each length $T$.
@@ 3.3 Predictive comparison via simulation
We compare our new GNARI and NGNAR models with the recent PNAR count network time series model of Armillotta and Fokianos (2024) via simulation. Using the same network as earlier (as shown in Figure 1) we simulate $500$ realizations of length $T=500$ from each of the following processes
@@
For each simulated realisation, we fit the following models: (A) GNARI(1,[1]1delimited-[]11,[1]1 , [ 1 ]), (B) NGNAR(1,[1]1delimited-[]11,[1]1 , [ 1 ]) fitted by conditional least squares and (C) by conditional maximum likelihood, (D) PNAR(1,[1]1delimited-[]11,[1]1 , [ 1 ]).
@@
A 145.0 209.5 9.08 211.0
B 145.0 209.5 8.96 211.0
C 145.1 209.4 8.96 211.0
D 145.2 209.7 9.08 211.8
(c) $T=50$
Table 3: Average mean-squared prediction error (MSPE) between the predicted next $T$ days time series by each model (A, B, C, D) fitted to training realizations from simulated processes P1, P2, P3 and P4 on the training set evaluated on the test set over 500 realizations from each simulated model.
Table 3 shows that for the simulated GNARI, NGNAR with positive parameters and PNAR processes, all four methods have almost equal performance with the PNAR model perhaps slightly worse than the others. For the NGNAR simulation with $\beta_{1,1}=-0.8$, the NGNAR models definitively predict better than the other models, especially over the short term. However, for longer horizons there is not much to choose between the methods.
The NGNAR model is clearly more flexible than the GNARI model and the PNAR model as it can cope with negative parameters. In principle, the log-linear version of the Armillotta and Fokianos (2024) model can cope with negative parameters. However, the error structure that the log-linear model assumes is different to that of the simulated processes P1–P4 above. To avoid doubt we repeated the simulation/prediction exercise above for the log-linear model and the prediction errors were uniformly at least four times worse than those reported in Table 3 and, in many cases, much worse.
GOT HERE
@@ 4.1 Data description
We obtained daily counts of people who tested positive for COVID19 in the 62 counties of New York State, USA from New York State Department of Health (2022) during the period 1st March 2020 to 23rd May 2022. The counts can be written as a multivariate time series of dimension $T\times N=783\times 62$ (some common missing values at the head of the time series were discarded).
@@ 4.2 Prediction evaluation method
We will fit four model types on the training series: GNAR, PNAR, GNARI and NGNAR models. Each model will be fitted twice using maximum lags of $14$ and $21$, respectively. The order of $\alpha$, $[I_{1},\dots,I_{p}]$ where $I_{j}$ is either 1 or 0 indicating whether the autoregressive term at lag $j$ is included. The order of $\beta$, $[s_{1},\ldots,s_{p}]$, will be selected using backward deletion and the Bayesian information criterion (BIC) as the metric. For the PNAR model, the quasi-MLE fitting method provided in the PNAR package in R can only fit a PNAR(1) model for this particular dataset as higher order models will lead to a non-zero score function.
We will fit the GNAR model using conditional least squares, the GNARI model using constrained least squares to ensure that the model parameters are non-negative. The NGNAR models will be fitted using conditional MLE using the ADAM optimizer from Kingma and Ba (2015). In the GNARI model we will assume that the $\epsilon_{i,t}$ is Poisson-distributed. The NGNAR models will use the $\operatorname{softplus}$ function as the response function and Poisson as the conditional distribution.
@@ 4.3 Results
NGNAR($21$) 3.97 6.33 4.53 16.5 87.3 306
PNAR($1$) 3.95 7.91 6.44 18.4 87.0 297
Table 4: The mean-squared prediction error ×$\times 100$ (MSPE) between the predictions up to forecast horizon $h$ (to three significant figures).
@@ 6 Acknowledgements
Li gratefully acknowledges partial support by Imperial College London. Both gratefully acknowledge support from EPSRC NeST Programme grant EP/X002195/1.
@@ Appendix A Proofs
Hence, the GNARI process (17) can be viewed as a multivariate integer-valued autoregressive (MGINAR) process, see Latour (1997):
@@
Latour (1997) Proposition 3.1 permits us to conclude that, under the condition that all roots of $\det(I-Az)$ are outside the unit circle, or, equivalently, all eigenvalues of $A$ are inside the unit circle, the process $\{\mathbf{X}_{t}\}_{t}$ has a unique stationary solution.
@@
Knight et al. (2020) Appendix A tells us that the above condition can be achieved if condition (19) holds. Hence, condition (19) guarantees the unique stationary solution of the GNARI process. ∎
Proof of Lemma (2.3).
@@
where $A$ is the same matrix as in (48), and $g(x)=s(x)$ is the $\operatorname{softplus}$ function. Then $\{\mathbf{Y}_{t}\}_{t}$ is Markov, aperiodic and irreducible and we have
@@
where $f(\mathbf{X})=\{g(X_{1}),\ldots,g(X_{p}),X_{p+1},\ldots\}^{T}$. From Appendix A in Knight et al. (2020), we know that all eigenvalues of $A$ can be assumed to be inside in the unit circle. This is equivalent to the spectral radius of $A$, $\rho(A)<1$. Thus, by Lemma 2.5 in An and Huang (1996), there exists a matrix norm $||\cdot||_{m}$, a vector norm $||\cdot||_{v}$, and $\lambda\in(0,1)$, such that
$||Ax||_{v}\leq||A||_{m}||x||_{v}\leq\ \lambda||x||_{v},\ \ \ \forall x\in\mathbb{R}^{Np}.$ (52)
@@
$||A\mathbf{y}||_{v}\leq\ \lambda||\mathbf{y}||_{v}.$ (54)
It is simple to show that $\lim_{x\xrightarrow{}\infty}g(x)=x$, and $|g(x)|<|x|$ for all $x<\operatorname{arcsinh}(-1/2)$, for $g(x)=s_{1}(x)$, the $\operatorname{softplus}$ function. We can thus find $C\in\mathbb{R}^{Np}$ to be $R^{Np}_{\geq 0}\bigcap C^{\prime}$ where $C^{\prime}$ is a sphere in $\mathbb{R}^{Np}$ such that if $\mathbf{y}\notin C$, the negative elements of $A\mathbf{y}$ are less than $\text{archsinh}(-1/2)$ and $||\mathbf{y}^{\prime}-A\mathbf{y}||_{v}<(1-\lambda)M/2$, where $\mathbf{y}^{\prime}$ is the vector whose non-negative entries equals that of $f(A\mathbf{y})$ and negative entries equals that of $A\mathbf{y}$, and $M=\inf_{\mathbf{y}\notin C}||Y||_{v}$.
@@
$||f(A\mathbf{y})||_{v}$ $\leq||A\mathbf{y}^{\prime}||_{v}$
$\leq||\mathbf{y}^{\prime}-A\mathbf{y}||_{v}+||A\mathbf{y}||_{v}$
$\leq(1-\lambda)M/2+\lambda||\mathbf{y}||_{v}$
$\leq(1/2+\lambda/2)||\mathbf{y}||_{v}$
where clearly $11/2+\lambda/2<$ 1.
Now, by Lemma 2.2 from An and Huang (1996), which is a reformulation of the Tweedie’s criterion for ergodicity:

### v2

@@ 2.2.1 NGNAR model definition
NGNAR models that are restricted to only having stage one neighbours and lag one autoregression are examples of the broad class of nonlinear network autoregressions introduced by nna. However, NGNAR models have the ability to model associations using more general autoregressive lags, $p$, and $r$-stage neighbours, which have proved important and effective for good network time series modelling. The more general models require modelling tools to select model order, as regular ARIMA(p,d,q)(p,d,q) models do, such as AIC, BIC or network auto- and partial autocorrelations and visualizations of these such as Corbit plots, see corbit23.
@@
  • The exponential response: $g(x)=\exp(x)$: in which case the model is similar to the log-linear Poisson autoregression model in loglinear, but replacing the $X$ by $\log(X+1)$, to prevent explosion as noted on page 564 of loglinear.
@@
Our exposition below uses the $\operatorname{softplus}$ response function with $c=1$, i.e., $s_{c}(x)=s_{1}(x)$. Estimation can be performed either by conditional least squares or conditional maximum likelihood and both are discussed below.
@@ 2.2.2 Stationarity and ergodicity for NGNAR
2.2.3 Existence of the moments of NGNAR
Lemma 2.4.
Assuming that $D$ is the Poisson distribution and that $X_{i,t}|\mathcal{F}_{t-1}$ are mutually independent, then (38) is also a sufficient, but not necessary, condition for static-network NGNAR($p,[s_{1},\ldots,s_{p}]$) processes to have $\mathbb{E}[\prod_{i=1}^{N}X_{i,t}^{m_{i}}]<\infty$ for all $t\geq 0$, $m_{i}\geq 0$.
@@ 2.2.4 Remarks on conditional least squares estimation
Again, for notational simplicity, we consider the global $\alpha$ case. Let $\mathbf{\theta}=(\alpha_{1},\beta_{1,1},\dots,\beta_{1,s_{1}},\dots,\alpha_{p},\beta_{p,1},\dots,\beta_{p,s_{p}},\alpha_{0})^{T}$ be the parameter of interest, $\mathbf{y}$ be the target vector as defined in (28), and $X$ be the design matrix as defined in (29). Then, the conditional least squares estimator, $\hat{\mathbf{\theta}}$, is the one that minimizes $Y-\mathbf{g}(X\mathbf{\theta})||^{2}$, where $\mathbf{g}(\mathbf{x})=\{g(x_{1}),g(x_{2}),\ldots\}^{T}$.
As the process is no longer linear by design, we can not use the usual linear least squares estimation method. Instead, we can use numerical methods such as gradient descent or the ADAM optimization in adam. Choosing the solution to $X^{T}X\theta=X^{T}Y$ as the initial value can speed up the optimization.
Proposition 3.
Under the assumption of Lemma 2.4, we have
  1. 1.
    $\hat{\theta}^{(n)}\xrightarrow{a.s}\mathbf{\theta}_{0}.$ (39)
  2. 2.
    $n^{1/2}(\hat{\mathbf{\theta}}^{(n)}-\mathbf{\theta}_{0})\xrightarrow{}\operatorname{MVN}(0,U^{-1}RU^{-1}),$ (40)
    as $n\xrightarrow{}\infty$, where $\operatorname{MVN}$ is the multivariate normal distribution, $U$ and $R$ are analogous to that in Proposition 1.
2.2.5 Quasi-maximum likelihood estimation
Unlike the GNARI model, the NGNAR model explicitly defines the conditional distribution of $X_{i,t}|\mathcal{F}_{t-1}$. Thus, it is feasible to implement the quasi-maximum likelihood estimator (QMLE). The QMLE $\hat{\mathbf{\theta}}_{M}$ maximizes the quasi-likelihood
$L(\mathbf{\theta})=\prod_{t=p}^{n-1}\prod_{i=1}^{N}f(X_{i,t+1}|\mathcal{F}_{t},\mathbf{\theta}),$ (41)
where $f(\mathbf{X_{i,t+1}}|\mathcal{F}_{t},\mathbf{\theta})$ is the density function for $X_{i,t+1}|\mathcal{F}_{t}$. We also have that $\mathbb{E}[\mathbf{X_{i,t+1}}|\mathcal{F}_{t},\mathbf{\theta}]=g([X]_{t-p+i}\cdot\mathbf{\theta})$, where $X]_{r}$ is the $r^{th}$ row of matrix $X$ defined in (29). We can then recognize that the objective function under the NGNAR model assumption is of the same form as that of a generalized linear model (GLM). Thus, the solution for $\hat{\mathbf{\theta}}_{M}$ can be computed using similar techniques as used for GLM, such as iterative weighted least squares.
If we assumed that the conditional distribution was Poisson, then the parameters estimated by conditional least squares would differ from those estimated by QMLE unless $X_{i,t}$ is large enough so that the Poisson conditional distribution can be approximated with a Gaussian conditional distribution. Then the two estimation methods should give similar results. If $X_{i,t}|\mathcal{F}_{t-1}$ are mutually independent, the quasi-log-likelihood is
$l_{n}(\theta)=\sum_{t=p+1}^{n}\sum_{i=1}^{N}X_{i,t}\log(M_{i,t})-M_{i,t}-\log(X_{i,t}!)$ (42)
Proposition 4.
Under the assumption of Lemma 2.4, we have
  1. 1.
    $\hat{\theta}^{(n)}_{M}\xrightarrow{a.s}\mathbf{\theta}_{0}.$ (43)
  2. 2.
    $n^{1/2}(\hat{\mathbf{\theta}}^{(n)}-\mathbf{\theta}_{0})\xrightarrow{}\operatorname{MVN}(0,U^{-1}RU^{-1}),$ (44)
    where $U$ and $R$ are defined in the proof in the appendix.
@@ 3.1 GNARI parameter estimation simulation study
This section investigates estimation performance using conditional least squares for a Poisson-GNARI($1,[1]$) process with $\alpha_{1}=0.5$, $\beta_{1,1}=0.4$, and $\lambda=10$ on a $N=50$ chain network shown with its adjacency matrix in Figure 1.
$M=\begin{bmatrix}0&1&&&&1\\ 1&0&1&&&\\ &&\ddots&&&\\ &&&\ddots&&\\ &&&1&0&1\\ 1&&&&1&0\end{bmatrix}$
Figure 1: 50-node chain network for simulation experiments. Left: picture of the network. Right: adjacency matrix.
We will simulate realizations from the Poisson-GNARI($1,[1]$) process with lengths $T=10,50,200$ and $500$ observations and repeat this $1000$ times for each choice of $T$ and estimate parameters for each realization.
@@
10 0.494 (0.039) 0.392 (0.053) 11.48 (5.08)
50 0.497 (0.017) 0.397 (0.021) 10.61 (1.99)
200 0.500 (0.0080) 0.399 (0.010) 10.15 (0.93)
500 0.500 (0.0053) 0.400 (0.0070) 10.07 (0.63)
True 0.500 0.400 10.0
Table 1: The mean (and standard deviation) over $1000$ conditional least squares estimates of each parameter in the GNARI($1,[1]$) model for each length $T$
Table 1 shows the results: the mean of the estimates clearly approaches the truth as $T$ gets larger. The standard deviation is approximately inversely proportional to $\sqrt{T}$, i,e., $\frac{\text{std}(\hat{\theta}_{T})}{\text{std}(\hat{\theta}_{m})}\approx\sqrt{\frac{m}{T}}$, which is consistent with the asymptotic properties. It is also worth noting that there is a tendency for underestimation of $\alpha_{1}$ and $\beta_{1,1}$, but overestimation of $\lambda$.
@@ 3.2 NGNAR parameter estimation simulation study
Table 2 shows the results of a similar simulation study to the previous one, but for a NGNAR process. We simulate 1000 realizations of a Poisson-NGNAR($1,[1]$) with $g(\cdot)=\operatorname{softplus}(\cdot)$, $\alpha_{1}=0.5$, $\beta_{1,1}=-0.4$, and $\alpha_{0}=10$ for lengths $T=10,50,200,500$ on the same network as in the previous section.
For each realization, we fit the NGNAR model using both conditional least squares and conditional maximum likelihood.
@@
500 CLS 0.500 -0.400 10.0
(0.0053) (0.0066) (0.115)
CMLE 0.500 -0.400 10.0
(0.0051) (0.0062) (0.112)
True 0.500 -0.400 10
Table 2: The mean (and standard deviation) across the 1000 conditional least squares and conditional MLE estimates for each parameter in the Poisson-NGNAR($1,[1]$) model for each length $T$.
@@ 3.3 Predictive comparison via simulation
We compare our new GNARI and NGNAR models with the recent PNAR count network time series model of pnar via simulation. Using the same network as earlier (as shown in Figure 1) we simulate $500$ realizations of length $T=500$ from each of the following processes
@@
For each simulated realisation, we fit the following models: (A) GNARI($1,[1]$), (B) NGNAR($1,[1]$) fitted by conditional least squares and (C) by conditional maximum likelihood, (D) PNAR($1,[1]$).
@@
A 129.8 192.6 10.31 194.5
B 129.8 192.6 10.07 194.5
C 129.8 192.6 10.07 194.5
D 129.8 192.6 10.31 194.5
(c) $h=50$
Table 3: Average mean-squared prediction error (MSPE) between the predicted next $h$ days time series by each model (A, B, C, D) fitted to training realizations from simulated processes P1, P2, P3 and P4 on the training set evaluated on the test set over 500 realizations from each simulated model.
Table 3 shows that for the simulated GNARI, NGNAR with positive parameters and PNAR processes, all four methods have almost equal performance. For the NGNAR simulation with $\beta_{1,1}=-0.8$, the NGNAR models definitively predict better than the other models, especially over the short term. However, for longer horizons there is not much to choose between the methods.
The NGNAR model is clearly more flexible than the GNARI model and the PNAR model as it can cope with negative parameters. In principle, the log-linear version of the pnar model can cope with negative parameters. However, the error structure that the log-linear model assumes is different to that of the simulated processes P1–P4 above. To avoid doubt we repeated the simulation/prediction exercise above for the log-linear model and the prediction errors were uniformly at least four times worse than those reported in Table 3 and, in many cases, much worse.
@@ 4.1 Data description
We obtained daily counts of people who tested positive for COVID19 in the 62 counties of New York State, USA from NYcovid during the period 1st March 2020 to 23rd May 2022. The counts can be written as a multivariate time series of dimension $T\times N=783\times 62$ (some common missing values at the head of the time series were discarded).
@@ 4.2 Prediction evaluation method
We will fit four model types on the training series: GNAR, PNAR, GNARI and NGNAR models. Each model will be fitted twice using maximum lags of $14$ and $21$, respectively. The order of $\alpha$, $I_{1},\dots,I_{p}]$ where $I_{j}$ is either 1 or 0 indicating whether the autoregressive term at lag $j$ is included. The order of $\beta$, $s_{1},\ldots,s_{p}]$, will be selected using backward deletion and the Bayesian information criterion (BIC) as the metric. For the PNAR model, the quasi-MLE fitting method provided in the PNAR package in R can only fit a PNAR(1) model for this particular dataset as higher order models will lead to a non-zero score function.
We will fit the GNAR model using conditional least squares, the GNARI model using constrained least squares to ensure that the model parameters are non-negative. The NGNAR models will be fitted using conditional MLE using the ADAM optimizer from adam. In the GNARI model we will assume that the $\epsilon_{i,t}$ is Poisson-distributed. The NGNAR models will use the $\operatorname{softplus}$ function as the response function and Poisson as the conditional distribution.
@@ 4.3 Results
NGNAR(2121) 3.97 6.33 4.53 16.5 87.3 306
PNAR(11) 3.95 7.91 6.44 18.4 87.0 297
Table 4: The mean-squared prediction error $\times 100$ (MSPE) between the predictions up to forecast horizon $h$ (to three significant figures).
@@ 6 Acknowledgements
Liu gratefully acknowledges partial support by Imperial College London. Both gratefully acknowledge support from EPSRC NeST Programme grant EP/X002195/1.
@@ Appendix A Proofs
Hence, the GNARI process (17) can be viewed as a multivariate integer-valued autoregressive (MGINAR) process, see mginar:
@@
mginar Proposition 3.1 permits us to conclude that, under the condition that all roots of $\det(I-Az)$ are outside the unit circle, or, equivalently, all eigenvalues of $A$ are inside the unit circle, the process $\{\mathbf{X}_{t}\}_{t}$ has a unique stationary solution. mginar Section 5 also shows that $\{\mathbf{X}_{t}\}_{t}$ is ergodic.
@@
gnar Appendix A tells us that the above condition can be achieved if condition (19) holds. Hence, condition (19) guarantees the unique stationary solution of the GNARI process. ∎
Proof of Proposition 1.
Assume that GNARI process is stationary. As in Section 2.1.5 define
$\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})=\mathbb{E}_{\mathbf{\theta}}[\mathbf{X_{t}}|\mathcal{F}_{t-1}].$ (58)
and let $\mathbf{\theta}_{0}$ be the true value of the (vector) parameter of interest $\mathbf{\theta}$. In the GNARI case, we have that
$\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})=\sum_{j=1}^{p}A_{j}\mathbf{X_{t-j}}+\lambda\mathbf{1}_{N}.$ (59)
Then
$\mathbb{E}\left[\left\|\frac{\partial\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})}{\partial\alpha_{j}}\right\|^{2}\right]=\mathbb{E}[\mathbf{X}_{t-j}^{T}\mathbf{X}_{t-j}]:=\eta_{\alpha}$ (60)
and
$\mathbb{E}\left[\left\|\frac{\partial\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})}{\partial\beta_{j,r}}\right\|^{2}\right]=\mathbb{E}[\mathbf{X}_{t-j}^{T}{W^{(r)}}^{T}W^{(r)}\mathbf{X}_{t-j}]:=\eta_{\beta},$ (61)
say, are expectations of quadratic forms in $\mathbf{X}_{t-j}$. By (22) and (23), we have both $\eta_{\alpha},\eta_{\beta}<\infty$. Further, $\mathbb{E}\left[\left\|\partial\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})/\partial\lambda\right\|^{2}\right]=\mathbf{1}^{T}_{N}\mathbf{1}_{N}=N<\infty$. We also know that all the second-order derivatives with respect to $\alpha_{j}$, $\beta_{j,r}$, or $\lambda$ are all zero, as $\hat{\mathbf{X}}_{t|t-1}$ depends linearly on each parameter.
Let $d$ be the length of the $\theta$ vector. Suppose there exists constants $a_{1},\dots,a_{d}$ such that
$\mathbb{E}\left[\left\|\sum_{k=1}^{d}a_{k}\frac{\partial\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})}{\partial\theta_{k}}\right\|^{2}\right]=0$ (62)
which implies that, for all $i=1,\dots,N$, there is a linear relationship between the $X_{q,t-j}$ given by
$\sum_{j=1}^{p}\sum_{q\in\mathcal{N}_{j}(i)}C_{i,j,q}X_{q,t-j}+a_{d}=0,\quad\quad a.s.$ (63)
where $\mathcal{N}_{j}(i)={i}\cup\mathcal{N}^{(1)}(i)\cup\dots\cup\mathcal{N}^{(s_{j})}(i)$, $C_{i,j,q}=c_{i,j,q}a_{k}$ for some $k$ and $c_{i,j,q}$ independent of $a_{1},\dots,a_{d}$. This implies that $C_{i,j,q}=0$, i.e., $a_{1}=\dots=a_{d}=0$. Then, by Therorem 3.1 in estnonlinear, we know that $\hat{\theta}^{(n)}\xrightarrow{a.s}\mathbf{\theta}_{0}$.
Now, let
$R=\mathbb{E}\left[\frac{\partial\hat{\mathbf{X}}_{t|t-1}^{T}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})f_{t|t-1}(\mathbf{\theta}_{0})\frac{\partial\hat{\mathbf{X}}_{t|t-1}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})\right],$ (64)
where as we assumed conditional independence,
$f_{t|t-1}(\mathbf{\theta})$ $=\mathbb{E}[\{\mathbf{X_{t}}-\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})\}\{\mathbf{X_{t}}-\hat{\mathbf{X}}_{t|t-1}(\mathbf{\theta})\}^{T}|\mathcal{F}_{t-1}]$ (65)
$=\text{diag}(\{\textbf{Var}(X_{i,t}|\mathcal{F}_{t-1})\}_{i}),$ (66)
with $\textbf{Var}(X_{i,t}|\mathcal{F}_{t-1})$ as in (15). In order to prove asymptotic normality, we require that $\|R\|_{2}<\infty$. In our case, to temporarily simplify notation, let $\theta_{a}=\beta_{j_{1},r_{1}}$, $\theta_{b}=\beta_{j_{2},r_{2}}$, then
$R_{a,b}=\mathbb{E}\left[\sum_{i=1}^{N}\{W^{(r_{1})}X_{t-j_{1}}\}_{i}\textbf{Var}(X_{i,t}|\mathcal{F}_{t-1})\{W^{(r_{2})}X_{t-j_{2}}\}_{i}\right]$ (67)
and similarly for the other entries. Since we have assumed that the $\epsilon_{i,t}$ are Poisson distributed, we have $\mathbb{E}[|\epsilon_{i,t}|^{3}]<\infty$, which implies that $\mathbb{E}[|X_{i,t}|^{3}]<\infty$, see franke1993multivariate. We can then use the Cauchy-Schwarz inequality to show that $|R_{a,b}|<\infty$, for all possible pairs $a,b$.
Hence, estnonlinear Theorem 3.2 implies that as $n\xrightarrow{}\infty$ we have
$n^{1/2}\{\hat{\mathbf{\theta}}^{(n)}-\mathbf{\theta}_{0}\}\xrightarrow{}\operatorname{MVN}(0,U^{-1}RU^{-1}),$ (68)
where
$U=\mathbb{E}\left[\frac{\partial\hat{\mathbf{X}}_{t|t-1}^{T}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})\frac{\partial\hat{\mathbf{X}}_{t|t-1}}{\partial\mathbf{\theta}}(\mathbf{\theta}_{0})\right].$ (69)
Proof of Proposition 2.
The proof of estnonlinear Theorem 3.1 shows that the regularity conditions of Theorem 3.1 which we have checked in our case as above implies the conditions of Theorem 2.1. The conclusion that $\hat{\theta}^{(n)}\xrightarrow{a.s}\mathbf{\theta}_{0}$ is a direct result of Theorem 2.1. We now show that Theorem 2.1 holds if we replace $\hat{\theta}^{(n)}$ with $\hat{\theta}^{(n)}_{[0,1]}$.
First note that $Q_{n}(\theta)$ is globally convex in our case, so we do not need to consider local minima. The proof of estnonlinear Theorem 2.1 states that for any $\epsilon,\delta>0$, there exists an event $E$ with $P(E)>1-\epsilon$ and $n_{0}\in\mathbb{N}$ such that on $E$, for any $n>n_{0}$ and $\theta$ on the boundary of $B_{\delta^{*}}(\theta^{0})$ where $0<\delta^{*}<\delta$ and $B_{\delta^{*}}(\theta^{0})$ is the open sphere of radius $\delta^{*}$ centered at $\theta^{0}$,
$Q_{n}(\theta)\geq Q_{n}(\theta^{0})$ (70)
Moreover, the minimum $\hat{\theta}^{(n)}$ is in $B_{\delta^{*}}(\theta^{0})$. This implies that $\hat{\theta}^{(n)}_{[0,1]}$ is in $B_{\delta^{*}}(\theta^{0})\cap[0,1]^{d}$ on event $E$. The rest is identical to the proof of Corollary 2.1 klimko1978. ∎
Proof of Lemma 2.3.
For the convenience of notation, we assume that $\alpha_{i,0}=0$ for all $i$.
@@
where $A$ is the same matrix as in (56), and $g(x)=s(x)$ is the entry-wise $\operatorname{softplus}$ function. Then $\{\mathbf{Y}_{t}\}_{t}$ is Markov, aperiodic and irreducible and we have
@@
where $f(\mathbf{X})=\{\underbrace{g(\mathbf{X}_{1})^{T}}_{N},\ldots,g(\mathbf{X}_{p})^{T},\mathbf{X}_{p+1}^{T},\ldots\}^{T}$. From Appendix A in gnar, we know that (38) implies that all eigenvalues of $A$ are inside the unit circle. This is equivalent to the spectral radius of $A$, $\rho(A)<1$. Thus, by Lemma 2.5 in ergodicitynlar, there exists a matrix norm $||\cdot||_{m}$, a vector norm $||\cdot||_{v}$, and $\lambda\in(0,1)$, such that
[…]
@@
[…]
@@
[…]
where clearly $1/2+\lambda/2<1$.
Now, by Lemma 2.2 from ergodicitynlar, which is a reformulation of the Tweedie’s criterion for ergodicity (Tweedie1975SufficientCF):